OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (0, 2) and (1, 5)?

OpenStudy (anonymous):

First, find the slope. Use the slope formula. \[m=\frac{ y_2-y_1 }{ x_2-x_1}\] Plug in the points. \[m=\frac{ 5-2 }{ 1-0 }\] The slope is m = 3. Plug it into the point slope formula along with one of the points. \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

3x -y + 2 = 0

OpenStudy (anonymous):

Right. That's a variation of y - 2 = 3(x - 0).